OpenStudy (thefluffmuncher):

1. Express the ratio in simplest form. 18 : 12 (Points : 4) 9 : 6 6 : 4 3 : 2 6 : 1

OpenStudy (thefluffmuncher):

pretty sure it's c but just making sure

OpenStudy (driftracer305):

yup....ur correct!!

OpenStudy (thefluffmuncher):

2. Express the ratio in simplest form. The ratio of 500 m to 2 km (Points : 4) 250 : 1 1 : 250 5 : 2 1 : 4

OpenStudy (thefluffmuncher):

not sure on this one

OpenStudy (thefluffmuncher):

but if i have to guess i go with a

OpenStudy (driftracer305):

k convert 2 km to m

OpenStudy (thefluffmuncher):

2000 m

OpenStudy (driftracer305):

right so now u have 500 : 2000

OpenStudy (thefluffmuncher):

so it's d

OpenStudy (driftracer305):

yes sir!!

OpenStudy (thefluffmuncher):

3. Solve:7/9=21/x (Points : 4) 3 14 27 14/9

OpenStudy (thefluffmuncher):

i'm going with c on this one

OpenStudy (driftracer305):

correct

OpenStudy (thefluffmuncher):

4. What is 30% of 60? (Points : 4) 1800 0.18 180 18

OpenStudy (thefluffmuncher):

it's d?

OpenStudy (driftracer305):

damn bro...u knw the answers..so y u asking?

OpenStudy (thefluffmuncher):

idk but i find comfort when someelse verifies stuff for even though i know what the anwser

OpenStudy (thefluffmuncher):

someone*

OpenStudy (thefluffmuncher):

5. Sixty is 5% of what number? (Points : 4) 3 12 300 1200

OpenStudy (thefluffmuncher):

on this one not sure how to do it

OpenStudy (thefluffmuncher):

@driftracer305

OpenStudy (thefluffmuncher):

@jdoe0001

OpenStudy (thefluffmuncher):

@girlover

sammixboo (sammixboo):

For the problem you just asked, you can use this formula :P \(\rm\Large \frac{part}{whole}=\frac{percent}{100}\) The part is 60, since we are finding out 60 is 5% of what number? The whole is unknown, but the number we are looking for The percent is 5%, so now we plug in our info! \(\rm\Large \frac{60}{whole}=\frac{5}{100}\) Now we can cross multiply, so we can multiply the numbers 60 and 100 together.